OpenStudy (anonymous):

@precal Write the answer in simplified radical form. assume that all variables represent real numbers.

OpenStudy (anonymous):

\[(2\sqrt{12}+7)(2\sqrt{12}-7)\]

OpenStudy (anonymous):

i attempted to solve it and got \[8\sqrt{6}-49\] the book says the answer is -1

OpenStudy (precal):